OpenStudy (anonymous):

The sum of two consecutive odd integers is at least 80. n + n + 2 ≥ 80 n + n + 2 > 80 n + n + 1 ≥ 80 n + n + 1 > 80

OpenStudy (anonymous):

bottom because if you remove the +1 you could put > or equal to

OpenStudy (anonymous):

but surely it should be n+(n+2) >= 80 because, if you have n to be odd, then n+1 will be even so it definitely needs to be n and n+2 otherwise they won't be both odd then to be 'at least' 80, means it can equal or be bigger than So I would have said the top one. But maybe i'm interpreting it incorrectly, maybe wait for a second opinion.

OpenStudy (anonymous):

i actuallt got the answer to this there is another question i have a question about

OpenStudy (anonymous):

What is the numerator of the fraction equivalent to 4/15* that has a denominator of 60?

OpenStudy (anonymous):

sara your right i didnt see that it was consecutive #'s

OpenStudy (anonymous):

okay so it means \[\frac{ 4 }{ 15 }= \frac{ ? }{ 60 }\] So if you cross multiply \[\frac{ 60*4 }{ 15 }=?\]

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

You are very welcome